# Hybrid Meta-learners for Estimating Heterogeneous Treatment Effects

### v1

@@ 6 Discussion
Meta-learning approaches remain a prominent strategy for estimating CATE due to their intuitive design and empirical performance. However, achieving robustness across diverse DGPs requires a deeper understanding of how different meta-learners encode inductive biases through their regularization strategies. Indirect learners regularize models for POs, while direct learners directly regularize the CATE function by constructing pseudo-outcome estimators. These differing inductive biases make certain strategies more effective under specific DGPs while failing on others.
H-learner introduces a generalized regularization strategy that interpolates between indirect and direct regularizations while adaptively tuning the appropriate strength of each. This adaptivity enables H-learner to dynamically adjust its inductive bias to better match the underlying structure of the data without requiring prior knowledge of the underlying DGPs. Experiments show that H-learner effectively overcomes the trade-off between indirect and direct learners, achieving robust performance by leveraging both regularization strategies.
Despite improved performance, our study has several key limitations. One limitation of H-learner is its increased computational cost, as selecting the optimal regularization level requires training multiple models across different $\lambda$ values and selecting the best based on validation performance. Additionally, our experiments are currently limited to neural network implementations, extending H-learner to other machine learning models is left as future work.
@@ Appendix A Implementations Details
A.1 Model Architectures and Training Details
We implement all models following the architecture used in [2]. Specifically, the representation network of H-learner consists of three fully connected layers, each with 200 units and exponential linear unit (ELU) activations. Each output head comprises two additional dense layers with 100 units, followed by a final prediction layer. We implement similar architectures for the other meta-learners to ensure a fair comparison. All models are trained for 1000 epochs using the AdamW optimizer with mini-batches of size 100 and a cosine annealing learning rate schedule. We perform a grid search over learning rates {0.0001, 0.0005, 0.001}, and checkpoint achieving the lowest validation loss is used for final evaluation. We detail the validation criteria in the following section.
While we propose the H-learner with sample splitting using $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ for first and second stages as is commonly done in the literature, we found that using the full dataset for both stages often yields better empirical performance, particularly in small-sample settings. Similar observations have been made in [18], where sample splitting is then omitted. We therefore do not apply sample splitting and instead train both stages using the full dataset $\mathcal{D}$. All experiments were conducted on a machine with an AMD EPYC 7543 32-Core Processor.
A.2 Hyperparameter Tuning
@@
To address this issue, we instead estimate the potential outcomes by training outcome models on the validation set, yielding $\check{\mu}_{0}(X_{i})$ and $\check{\mu}_{1}(X_{i})$. We then select the value of $\lambda$ that minimizes the following validation loss: $\frac{1}{n}\sum_{i=1}^{n}\left[\left(f_{1}(x_{i})-f_{0}(x_{i})\right)-\left(t_{i}(y_{i}-\check{\mu}_{0}(x_{i}))+(1-t_{i})(\check{\mu}_{1}(x_{i})-y_{i})\right)\right]^{2}.$ Similar strategies have also been employed in prior work [25, 26]. Note that all validations are performed using the same validation set as all other meta-learners, ensuring that H-learner uses the same amount of information for a fair comparison.
Appendix B Additional Semi-Synthetic Experiment Results
@@
Lastly, we add additional results by expanding Figure 5 to include all three experimental setups. Figure 7 presents the H-learner’s performance across different values of the regularization strength $\lambda$ for each setup. Note that $\lambda=0$ corresponds to the indirect learner, and $\lambda=1$ to the direct learner. Depending on the levels of feature sharing, treatment proportion, and confounding, either $\lambda=0$ or $\lambda=1$ may perform better. However, the optimal value of $\lambda$ consistently falls between the two extremes, demonstrating the strength of the H-learner in leveraging both types of regularization. Across different setups, the H-learner consistently outperforms both the direct and indirect learners.
Appendix C Additional Results on ACIC 2016
The ACIC 2016 benchmark datasets were introduced as part of the Atlantic Causal Inference Competition [24] and offer a comprehensive testbed for evaluating causal inference methods. The benchmark includes 77 distinct DGPs that vary in the complexity of the response surface, the degree of confounding, covariate overlap, and treatment effect heterogeneity. While the main paper reports results averaged over all DGPs, here we provide aggregated results for specific settings.
Results. We present the results in Table 3. The H-learner consistently achieves state-of-the-art performance across various response models, levels of heterogeneity, confounding, and overlap. The best performance is obtained when constructing the X-learner pseudo-outcome in the first stage. We also observe that using the DR-learner pseudo-outcome yields competitive performance under settings with low confounding or low heterogeneity. This aligns with expectations, as the DR pseudo-outcome estimator tends to suffer less from high variance in such scenarios.
Table 3: Within-sample and out-of-sample performance on the ACIC 2016 benchmark dataset across different settings.

### v2

@@ 7 Discussion
Meta-learning approaches remain a prominent strategy for estimating CATE due to their intuitive design and strong empirical performance. However, achieving robustness across diverse DGPs requires a deeper understanding of how different meta-learners encode inductive biases through their regularization strategies. Indirect learners regularize models for POs, while direct learners directly regularize the CATE function by constructing pseudo-outcome estimators. These differing inductive biases make certain strategies more effective under specific DGPs while failing on others.
H-learner introduces a generalized regularization strategy that interpolates between indirect and direct regularizations while adaptively tuning the appropriate strength of each. This adaptivity enables H-learner to adjust its inductive bias to better match the underlying structure of the data without requiring prior knowledge of the DGPs. Experiments show that H-learner effectively overcomes the trade-off between indirect and direct learners, and achieving robust performance.
H-learner also offers several promising directions for future extensions. Potential directions include integrating H-learner with more complex causal models, such as those involving hidden confounding or multi-level treatments. Further, our experiments are currently limited to neural network implementations, extending H-learner to other ML models is left for future work.
Acknowledgements
This work was supported by a Hellman Fellowship award from the Society of Hellman Fellows and the Evergreen Foundation.
@@ Appendix A Implementations Details
A.1 Pseudo-algorithm of H-learner
Below, we present the pseudo-code for the H-learner.
1
Input :   $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$
Output :   $\widehat{\theta}(x)=\widehat{f}_{1}(x)-\widehat{f}_{0}(x)$
2
3Split $\mathcal{D}$ into $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ ;
4
5Stage 1: Pseudo-outcome construction ;
7 Construct ${Y}_{\varphi}$ via the options in Table 1 ;
8
9Stage 2: Training the H-learner ;
11
return $\widehat{\theta}(x)=\widehat{f}_{1}(x)-\widehat{f}_{0}(x)$
Algorithm 1 H-learner
A.2 Model Architectures and Training Details
We implement all models following the architecture used in (Shalit et al., 2017). Specifically, the representation network of H-learner consists of three fully connected layers, each with 200 units and exponential linear unit (ELU) activations. Each output head comprises two additional dense layers with 100 units, followed by a final prediction layer. We implement similar architectures for the other meta-learners to ensure a fair comparison. All models are trained for 1000 epochs using the AdamW optimizer with mini-batches of size 100 and a cosine annealing learning rate schedule. We perform a grid search over learning rates {0.0001, 0.0005, 0.001}, and checkpoint achieving the lowest validation loss is used for final evaluation. We detail the validation criteria in the following section.
While we propose the H-learner with sample splitting using $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ for first and second stages as is commonly done in the literature, we found that using the full dataset for both stages often yields better empirical performance, particularly in small-sample settings. Similar observations have been made in Curth and Van der Schaar (2021), where sample splitting is then omitted. We therefore do not apply sample splitting and instead train both stages using the full dataset $\mathcal{D}$. Experiments were conducted on machine with an AMD EPYC 7543 32-Core Processor.
A.3 Hyperparameter Tuning
@@
To address this issue, we instead estimate the POs by training outcome models on the validation set, yielding $\check{\mu}_{0}(X_{i})$ and $\check{\mu}_{1}(X_{i})$. We then select the value of $\lambda$ that minimizes the following validation loss: $\frac{1}{n}\sum_{i=1}^{n}\left[\left(f_{1}(x_{i})-f_{0}(x_{i})\right)-\left(t_{i}(y_{i}-\check{\mu}_{0}(x_{i}))+(1-t_{i})(\check{\mu}_{1}(x_{i})-y_{i})\right)\right]^{2}.$ Similar strategies have also been employed in prior works (Alaa and Van Der Schaar, 2019; Mahajan et al., 2022). We compare the $\lambda$ chosen by our validation-based strategy with the ground-truth optimal $\lambda^{\ast}$ (minimizing test PEHE) in Table 3, showing that the validation-selected $\lambda$ closely tracks the optimal value. Note that all validations are performed using the same validation set as other meta-learners, ensuring H-learner uses the same amount of information for fair comparisons.
Feature Overlap Ratio $\lambda^{\ast}$ (Test PEHE) $\lambda$ (Validation X-score)
0.1 0.49 0.52
0.3 0.54 0.57
0.5 0.57 0.58
0.7 0.57 0.58
0.9 0.67 0.62
Table 3: Comparison between the optimal $\lambda^{\ast}$ minimizing test PEHE and the $\lambda$ selected using the X-score on the validation set in the first synthetic setup of Section 6.2.
Appendix B Proof and Additional Theoretical Analysis
B.1 Complete Proof for the H-Learner Solution in Linear Models
Setup. Let $X\in\mathbb{R}^{n\times d}$ denote the covariate matrix, $T\in\{0,1\}^{n}$ denote the treatment vector, and $Y\in\mathbb{R}^{n}$ denote the observed outcomes. Let $X_{1}\in\mathbb{R}^{n_{1}\times d}$ and $X_{0}\in\mathbb{R}^{n_{0}\times d}$ denote the submatrices of $X$ corresponding to treated ($T=1$) and control ($T=0$) units, and similarly $Y_{1}\in\mathbb{R}^{n_{1}}$ and $Y_{0}\in\mathbb{R}^{n_{0}}$. Let $Y_{\phi}\in\mathbb{R}^{n}$ be the pseudo-outcomes. For ease of notation, let $G_{1}=X_{1}^{\top}X_{1},\ G_{0}=X_{0}^{\top}X_{0},\ G=X^{\top}X,\ S_{1}=X_{1}^{\top}Y_{1},\ S_{0}=X_{0}^{\top}Y_{0},\ S_{\phi}=X^{\top}Y_{\phi}$.
Indirect and direct estimators. The indirect learner fits separate linear models to both groups:
$\hat{\theta}^{(1)}_{\mathrm{ind}}=(X_{1}^{\top}X_{1})^{-1}X_{1}^{\top}Y_{1}=G_{1}^{-1}S_{1},\qquad\hat{\theta}^{(0)}_{\mathrm{ind}}=(X_{0}^{\top}X_{0})^{-1}X_{0}^{\top}Y_{0}=G_{0}^{-1}S_{0},\qquad\hat{\theta}_{\mathrm{ind}}=\hat{\theta}^{(1)}_{\mathrm{ind}}-\hat{\theta}^{(0)}_{\mathrm{ind}}$
The direct estimator from the pseudo-outcome regression is
$\hat{\theta}_{\mathrm{dir}}=(X^{\top}X)^{-1}X^{\top}Y_{\phi}=G^{-1}S_{\phi}$
H-learner. We study the H-learner setup by modeling $f_{1}$ and $f_{0}$ as linear models with parameters $\theta_{1},\theta_{0}\in\mathbb{R}^{d}$:
$f_{1}(x)=\theta_{1}^{\top}x,\qquad f_{0}(x)=\theta_{0}^{\top}x,\qquad f_{1}(x)-f_{0}(x)=(\theta_{1}-\theta_{0})^{\top}x=\hat{\theta}_{H}^{\top}x$
We minimize the H-learner objective
$L(\theta_{1},\theta_{0})=(1-\lambda)\big(\|Y_{1}-X_{1}\theta_{1}\|_{2}^{2}+\|Y_{0}-X_{0}\theta_{0}\|_{2}^{2}\big)+\lambda\,\|X(\theta_{1}-\theta_{0})-Y_{\phi}\|_{2}^{2}$
Taking gradients and setting to zero yields
$\frac{\partial L}{\partial\theta_{1}}=-2(1-\lambda)X_{1}^{\top}(Y_{1}-X_{1}\theta_{1})+2\lambda X^{\top}(X(\theta_{1}-\theta_{0})-Y_{\phi})=0$
$\frac{\partial L}{\partial\theta_{0}}=-2(1-\lambda)X_{0}^{\top}(Y_{0}-X_{0}\theta_{0})-2\lambda X^{\top}(X(\theta_{1}-\theta_{0})-Y_{\phi})=0$
Rearrange:
$(1-\lambda)G_{1}\theta_{1}+\lambda G(\theta_{1}-\theta_{0})$ $=(1-\lambda)S_{1}+\lambda S_{\phi}$ (1)
$(1-\lambda)G_{0}\theta_{0}-\lambda G(\theta_{1}-\theta_{0})$ $=(1-\lambda)S_{0}-\lambda S_{\phi}$ (2)
Multiply (1) by $G_{1}^{-1}$ and (2) by $G_{0}^{-1}$:
$(1-\lambda)\theta_{1}+\lambda G_{1}^{-1}G(\theta_{1}-\theta_{0})$ $=(1-\lambda)G_{1}^{-1}S_{1}+\lambda G_{1}^{-1}S_{\phi},$
$(1-\lambda)\theta_{0}-\lambda G_{0}^{-1}G(\theta_{1}-\theta_{0})$ $=(1-\lambda)G_{0}^{-1}S_{0}-\lambda G_{0}^{-1}S_{\phi}$
Subtract the second from the first:
$\big[(1-\lambda)I+\lambda(G_{1}^{-1}+G_{0}^{-1})G\big](\theta_{1}-\theta_{0})=(1-\lambda)(G_{1}^{-1}S_{1}-G_{0}^{-1}S_{0})+\lambda(G_{1}^{-1}+G_{0}^{-1})S_{\phi}$
Then
$\big[(1-\lambda)I+\lambda A\big]\hat{\theta}_{H}=(1-\lambda)\hat{\theta}_{\text{ind}}+\lambda A\hat{\theta}_{\text{dir}},\quad\text{where }A=(G_{1}^{-1}+G_{0}^{-1})G$
Hence,
$\hat{\theta}_{H}$ $=\big[(1-\lambda)I+\lambda A\big]^{-1}\big[(1-\lambda)\hat{\theta}_{\text{ind}}+\lambda A\hat{\theta}_{\text{dir}}\big]$
$=(1-\lambda)\big[(1-\lambda)I+\lambda A\big]^{-1}\hat{\theta}_{\text{ind}}+\lambda A\big[(1-\lambda)I+\lambda A\big]^{-1}\hat{\theta}_{\text{dir}}$
Therefore,
$\hat{\theta}_{H}=(I-W)\hat{\theta}_{\text{ind}}+W\hat{\theta}_{\text{dir}},\quad W=\lambda A\big[(1-\lambda)I+\lambda A\big]^{-1}$
B.2 Bias–Variance Analysis of the H-Learner Estimator
B.2.1 Equivalence Between Parameter-Space MSE and CATE Prediction Error
Assumption (Isotropy).
Throughout this analysis, we assume the covariates are isotropic, i.e.,
$\Sigma_{X}:=\mathbb{E}[XX^{\top}]=cI_{d},\text{ for some constant }c>0.$
Case 1: When the true CATE function is linear $\tau(x)=x^{\top}\theta^{\star}$, then for any estimator $\hat{\theta}$, the CATE prediction error is
$\mathbb{E}\!\left[(\hat{\tau}(X)-\tau(X))^{2}\right]$ $=\mathbb{E}\!\left[(X^{\top}\hat{\theta}-X^{\top}\theta^{\star})^{2}\right]=\mathbb{E}\!\left[(X^{\top}(\hat{\theta}-\theta^{\star}))^{2}\right]$
$=(\hat{\theta}-\theta^{\star})^{\top}\,\mathbb{E}[XX^{\top}]\,(\hat{\theta}-\theta^{\star})=(\hat{\theta}-\theta^{\star})^{\top}\,\Sigma_{X}\,(\hat{\theta}-\theta^{\star})$ (3)
Under isotropy, $\Sigma_{X}=cI_{d}$, so (3) becomes
$\mathbb{E}\!\left[(\hat{\tau}(X)-\tau(X))^{2}\right]=c\,\|\hat{\theta}-\theta^{\star}\|_{2}^{2}.$
Therefore, comparing CATE error is equivalent (up to the constant) to comparing their parameter-space errors:
$\mathbb{E}\!\left[(\hat{\tau}(X)-\tau(X))^{2}\right]\;\propto\;\mathbb{E}\!\left[\|\hat{\theta}-\theta^{\star}\|_{2}^{2}\right]$
Case 2: When the true CATE function $\tau(x)$ is not linear, let
$\theta^{\star}:=\arg\min_{\theta}\ \mathbb{E}\big[(x^{\top}\theta-\tau(x))^{2}\big]$
be its best linear projection. Then for any estimator $\hat{\theta}$,
$\mathbb{E}\big[(x^{\top}\hat{\theta}-\tau(x))^{2}\big]$ $=\mathbb{E}\big[(x^{\top}\theta^{\star}-\tau(x))^{2}\big]+\mathbb{E}\big[(x^{\top}\hat{\theta}-x^{\top}\theta^{\star})^{2}\big]$
$=\underbrace{\mathbb{E}\big[(x^{\top}\theta^{\star}-\tau(x))^{2}\big]}_{\text{approximation error (independent of }\hat{\theta}\text{)}}\;+\;c\,\mathbb{E}\big[\|\hat{\theta}-\theta^{\star}\|_{2}^{2}\big].$
Therefore, even in the nonlinear case, all estimator comparisons can be carried out entirely in parameter space via $\mathbb{E}\|\hat{\theta}-\theta^{\star}\|_{2}^{2}$. Hence, in the following sections, we compare estimators through their parameter-space MSE.
B.2.2 Mean Squared Error Decomposition for the H-Learner
Let $e_{\mathrm{ind}}=\hat{\theta}_{\mathrm{ind}}-\theta^{\star},e_{\mathrm{dir}}=\hat{\theta}_{\mathrm{dir}}-\theta^{\star}$ denote the estimation errors of the indirect and direct estimators, denote the estimation errors of the indirect and direct estimators, where $\theta^{\star}\in\mathbb{R}^{d}$ is the true treatment-effect parameter in the linear CATE case, or the best linear projection as defined above in the nonlinear case.
Define their biases and covariances as
$b_{\mathrm{ind}}=\mathbb{E}[e_{\mathrm{ind}}],\qquad b_{\mathrm{dir}}=\mathbb{E}[e_{\mathrm{dir}}],\qquad\Sigma_{\mathrm{ind}}=\operatorname{Var}(e_{\mathrm{ind}}),\qquad\Sigma_{\mathrm{dir}}=\operatorname{Var}(e_{\mathrm{dir}})$
Then their mean squared errors can be written as
$\mathrm{MSE}_{\mathrm{ind}}=\mathbb{E}\big[\|\hat{\theta}_{\mathrm{ind}}-\theta^{\star}\|_{2}^{2}\big]=\|b_{\mathrm{ind}}\|^{2}+\operatorname{tr}(\Sigma_{\mathrm{ind}}),\qquad\mathrm{MSE}_{\mathrm{dir}}=\mathbb{E}\big[\|\hat{\theta}_{\mathrm{dir}}-\theta^{\star}\|_{2}^{2}\big]=\|b_{\mathrm{dir}}\|^{2}+\operatorname{tr}(\Sigma_{\mathrm{dir}})$
H-learner. We begin by analyzing the MSE along the scalar path $W=\omega I$, under which $\hat{\theta}_{H}$ forms a convex combination of the indirect and direct OLS estimators: $\hat{\theta}_{H}=(1-\omega)\,\hat{\theta}_{\mathrm{ind}}+\omega\,\hat{\theta}_{\mathrm{dir}}.$
We then compute the bias of the H-learner.
$b_{H}$ $=\mathbb{E}[\hat{\theta}_{H}-\theta^{\star}]$
$=\mathbb{E}\left[(1-\omega)\,\hat{\theta}_{\mathrm{ind}}+\omega\,\hat{\theta}_{\mathrm{dir}}-\theta^{\star}\right]$
$=(1-\omega)\,b_{\mathrm{ind}}+\omega\,b_{\mathrm{dir}}.$
Variance of the H-learner:
$\operatorname{Var}(\hat{\theta}_{H})$ $=\operatorname{Var}\big((1-\omega)\,\hat{\theta}_{\mathrm{ind}}+\omega\,\hat{\theta}_{\mathrm{dir}}\big)$
$=(1-\omega)^{2}\operatorname{Var}(\hat{\theta}_{\mathrm{ind}})+\omega^{2}\operatorname{Var}(\hat{\theta}_{\mathrm{dir}})+2\omega(1-\omega)\,\operatorname{Cov}(\hat{\theta}_{\mathrm{ind}},\hat{\theta}_{\mathrm{dir}}).$
Therefore, the MSE of the H-learner is:
$\mathrm{MSE}_{H}$ $=\mathbb{E}\!\left[\big\|\hat{\theta}_{H}-\theta^{\star}\big\|_{2}^{2}\right]$
$=\|b_{H}\|_{2}^{2}+\operatorname{tr}\!\big(\operatorname{Var}(\hat{\theta}_{H})\big)$
$=\big\|(1-\omega)\,b_{\mathrm{ind}}+\omega\,b_{\mathrm{dir}}\big\|_{2}^{2}+\operatorname{tr}\!\Big((1-\omega)^{2}\Sigma_{\mathrm{ind}}+\omega^{2}\Sigma_{\mathrm{dir}}+2\omega(1-\omega)\Sigma_{\mathrm{ind,dir}}\Big)$
$=(1-\omega)^{2}\!\left(\|b_{\mathrm{ind}}\|_{2}^{2}+\operatorname{tr}(\Sigma_{\mathrm{ind}})\right)+\omega^{2}\!\left(\|b_{\mathrm{dir}}\|_{2}^{2}+\operatorname{tr}(\Sigma_{\mathrm{dir}})\right)+2\omega(1-\omega)\!\left(b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}+\operatorname{tr}(\Sigma_{\mathrm{ind,dir}})\right)$
$=(1-\omega)^{2}\,\mathrm{MSE}_{\mathrm{ind}}+\omega^{2}\,\mathrm{MSE}_{\mathrm{dir}}+2\omega(1-\omega)\,D,$
where $D=b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}+\operatorname{tr}\big(\Sigma_{\mathrm{ind,dir}}\big),\ \ \Sigma_{\mathrm{ind,dir}}=\operatorname{Cov}(\hat{\theta}_{\mathrm{ind}},\hat{\theta}_{\mathrm{dir}})$
B.2.3 When Does the H-Learner Strictly Improve MSE?
We now analyze conditions that the optimal convex combination $\hat{\theta}_{H}=(1-\omega)\,\hat{\theta}_{\mathrm{ind}}+\omega\,\hat{\theta}_{\mathrm{dir}}$ achieve strictly smaller MSE than either endpoint ($\omega=0$ or $\omega=1$).
From the quadratic form of the H-learner MSE, the optimal weight is
$\omega^{\star}=\frac{\mathrm{MSE}_{\mathrm{ind}}-D}{\mathrm{MSE}_{\mathrm{ind}}+\mathrm{MSE}_{\mathrm{dir}}-2D},\qquad$
The H-learner strictly improves on both endpoints if and only if $\omega^{\star}\in(0,1)$, which holds when
$D<\min\{\mathrm{MSE}_{\mathrm{ind}},\,\mathrm{MSE}_{\mathrm{dir}}\}$ (4)
Now we prove Corollary 5.4 and 5.5:
Under the assumption of cross-fitting, $\Sigma_{\mathrm{ind,dir}}=\operatorname{Cov}(\hat{\theta}_{\mathrm{ind}},\hat{\theta}_{\mathrm{dir}})=0,D=b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}$.
  • Case 1: If $b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}<0$, then $D<0\leq\min\{\mathrm{MSE}_{\mathrm{ind}},\mathrm{MSE}_{\mathrm{dir}}\}$, so condition (4) holds.
  • Case 2: If $\|b_{\mathrm{ind}}\|_{2}\geq\|b_{\mathrm{dir}}\|_{2}$ and $\operatorname{Var}(\hat{\theta}_{\mathrm{ind}})=\operatorname{tr}(\Sigma_{\mathrm{dir}})>b_{\mathrm{dir}}^{\top}(b_{\mathrm{ind}}-b_{\mathrm{dir}})$. Then
    $D=b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}\leq\|b_{\mathrm{ind}}\|_{2}\,\|b_{\mathrm{dir}}\|_{2}\leq\|b_{\mathrm{ind}}\|_{2}^{2}<\|b_{\mathrm{ind}}\|_{2}^{2}+\operatorname{tr}(\Sigma_{\mathrm{ind}})=\mathrm{MSE}_{\mathrm{ind}}$
    $D=b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}<b_{\mathrm{dir}}^{\top}b_{\mathrm{dir}}+\operatorname{tr}(\Sigma_{\mathrm{dir}})=\mathrm{MSE}_{\mathrm{dir}}$
    Therefore $D<\min\{\mathrm{MSE}_{\mathrm{ind}},\mathrm{MSE}_{\mathrm{dir}}\}$, so condition (4) holds.
Since $\min_{W}\mathrm{MSE}_{H}(W)\leq\min_{\omega\in[0,1]}\mathrm{MSE}_{H}(\omega I)$, the conditions derived above along this scalar path provide a sufficient guarantee: whenever the scalar path admits an interior minimizer, the full matrix optimization achieves no higher risk and therefore strictly improves upon both endpoints.
Appendix C Additional Semi-Synthetic Experiment Results
@@
Effect of regularization parameter $\lambda$. Lastly, we add additional results by expanding Figure 5 to include all three experimental setups. Figure 8 presents the H-learner’s performance across different values of the regularization strength $\lambda$ for each setup. Note that $\lambda=0$ corresponds to the indirect learner, and $\lambda=1$ to the direct learner. Depending on the levels of feature sharing, treatment proportion, and confounding, either $\lambda=0$ or $\lambda=1$ may perform better. However, the optimal value of $\lambda$ consistently falls between the two extremes, demonstrating the strength of the H-learner in leveraging both types of regularization. Across different setups, the H-learner consistently outperforms both the direct and indirect learners.
@@
Appendix D Additional Results on ACIC 2016
The ACIC 2016 benchmark datasets were introduced as part of the Atlantic Causal Inference Competition (Dorie et al., 2019) and offer a comprehensive testbed for evaluating causal inference methods. The benchmark includes 77 distinct DGPs that vary in the complexity of the response surface, the degree of confounding, covariate overlap, and treatment effect heterogeneity. While the main paper reports results averaged over all DGPs, here we provide aggregated results for specific settings.
Results. We present the results in Tables 4, 5, 6, and 7. The H-learner consistently achieves state-of-the-art performance across various response models, levels of heterogeneity, confounding, and overlap. The best performance is obtained when constructing the X-learner pseudo-outcome in the first stage. We also observe that using the DR-learner pseudo-outcome yields competitive performance under settings with low confounding or low heterogeneity. This aligns with expectations, as the DR pseudo-outcome estimator tends to suffer less from high variance in such scenarios.
Table 4: Performance on the ACIC 2016 benchmark dataset across different response models.